The sides of a triangle are in the ratio 5: 12: 13. What is the length of each side of the triangle if the perimeter of the triangle is 15 in.?
Answer by mananth(16949)

5:12:13
The sides will be 5x,12x,13x
5x+12x+13x= 15
30x=15
x= 0.5
the sides are
0.5*5 = 2.5in
12*0.5=6in
13*0.5=6.5in